# SuperiorGAT: Graph Attention Networks for Sparse LiDAR Point Cloud Reconstruction in Autonomous Systems

### v2

@@ Introduction
Figure 5: Sensitivity analysis of the neighborhood size $k$ illustrating (a) the relationship between vertical reconstruction error ($RMSE_{z}$) and inference latency, and (b) the corresponding geometric shape fidelity measured by Chamfer distance.
Our results indicate that $k=10$ provides the peak performance, achieving the lowest vertical error with an $RMSE_{z}$ of 0.1808 m. While increasing $k$ beyond this point improves the global geometric shape (as shown by the decreasing Chamfer distance in Fig. 5b), it leads to a slight degradation in vertical precision ($RMSE_{z}$) due to over-smoothing of local terrain features. Furthermore, $k=10$ maintains a highly efficient inference latency of 8.63 s, demonstrating a favorable balance between computational demand and reconstruction accuracy. Based on these findings, $k=10$ was selected as the optimal configuration and is utilized for all subsequent experiments across diverse environments, including the KITTI campus, person, city, and road datasets.
The observed performance trends in Fig. 5 highlight the trade-off between local feature extraction and global context. When $k$ is small ($k=5$), the graph is too sparse to capture sufficient spatial relationships, leading to higher reconstruction errors. Conversely, at $k=20$, the inclusion of distant neighbors introduces noise and leads to a "smoothing" effect that obscures fine terrain details, which is reflected in the rising $RMSE_{z}$. The linear increase in latency from 7.08 s to 12.09 s further confirms that $k=10$ is the optimal "elbow point" for computational efficiency. Having established $k=10$ as the optimal hyperparameter, we now evaluate the robustness of SuperiorGAT by comparing it against state-of-the-art baselines across multiple diverse LiDAR environments.
Following the sensitivity analysis, we evaluate the SuperiorGAT model against traditional interpolation techniques and deep learning baselines. This comparison is conducted on the residential dataset using the optimized $k=10$ configuration. The quantitative results across four key metrics are illustrated in Fig.6.
Figure 6: Comparative benchmarking of SuperiorGAT ($k=10$) across: (a) Mean $RMSE_{z}$, (b) Mean $RMSE_{xyz}$, (c) Mean Chamfer Distance(m), and (d) Inference Latency(s).
The benchmarking results demonstrate that SuperiorGAT achieves superior global reconstruction accuracy. Specifically, it yields the lowest $RMSE_{xyz}$ (0.104 m) and Chamfer distance (0.143 m) among all tested methods. While Nearest Neighbor (NN) shows a slightly lower vertical error ($RMSE_{z}=0.174$ m compared to SuperiorGAT’s 0.181 m), this marginal gain is misleading in the context of 3D scene fidelity. NN interpolation is a purely local operator that preserves exact elevations but ignores spatial context, resulting in severe "staircase" artifacts and poor global coordinate alignment, as evidenced by its high $RMSE_{xyz}$ (0.330 m) and the second-worst Chamfer distance (0.546 m). SuperiorGAT effectively balances local vertical precision with global geometric consistency.
Regarding computational efficiency, SuperiorGAT exhibits a latency of 8.63 s, which is significantly faster than the standard GAT Baseline (13.08 s). While traditional methods, such as linear and NN interpolation, are near-instantaneous (<1<1 ms), they lack the learnable parameters required to model complex terrain geometries or adapt to beam dropout patterns. The increased latency in SuperiorGAT is a necessary trade-off for the attention mechanism’s ability to dynamically weight the importance of neighbors based on spatial features. While the current graph-based implementation reflects the computational intensity of neighbor searching, the refined SuperiorGAT architecture achieves a 34% reduction in processing time compared to deeper GAT baselines. This establishes a high accuracy-to-latency ratio for graph-based point cloud recovery and provides a viable foundation for future real-time optimization on specialized hardware.
To evaluate the robustness and adaptability of the proposed model across diverse operational environments, we conducted a comprehensive cross-dataset validation using five distinct KITTI sequences: Residential, City, Person, Campus, and Road. These environments encompass a wide spectrum of geometric complexities, ranging from structured urban corridors to dynamic scenes featuring pedestrians and open highway surfaces.
The comparative performance of SuperiorGAT against baseline and state-of-the-art methods is summarized in Table 1. The results demonstrate that SuperiorGAT consistently achieves superior reconstruction accuracy, maintaining an $RMSE_{xyz}$ of under 0.11m across all tested environments. Notably, while the Nearest Neighbor (NN) approach yields competitive $RMSE_{z}$ values in specific scenarios, it exhibits significantly higher Chamfer Distances—frequently exceeding double that of our model. This discrepancy indicates that NN fails to preserve the underlying structural geometry of the point cloud. In contrast, SuperiorGAT provides a balanced optimization between vertical precision and global spatial fidelity. Furthermore, our model maintains a consistent inference latency of approximately 8.6 s per frame, representing a significant computational improvement over standard multi-layer graph architectures and ensuring a high accuracy-to-latency ratio for complex spatial reconstruction.
@@
Dataset Method $RMSE_{z}$ (m) $RMSE_{xyz}$ (m) Chamfer Distance (m) Time (s)
@@
Figure 7 presents a qualitative comparison between the original ground truth LiDAR data and the reconstructed vertical profiles generated by SuperiorGAT across the Person, Road, Campus, and City datasets. Each panel provides a camera-view reference (left) for operational context alongside the corresponding 2D X-Z projection (right). For visual clarity, the reconstruction is shown using a sampling stride of 15, highlighting the model’s high geometric fidelity in recovering vertical structural height ($Z$) based on horizontal distance ($X$)
@@ Conclusion
This paper presented SuperiorGAT, a graph attention-based framework for reconstructing missing elevation information in sparse LiDAR point clouds under structured beam dropout. By modeling LiDAR scans as beam-aware graphs and enhancing standard graph attention networks with gated residual fusion and feed-forward refinement, the proposed approach effectively recovers lost vertical geometry while maintaining computational efficiency, without increasing network depth.
Experimental evaluations across diverse KITTI sequences, including Person, Road, Campus, and City environments, demonstrate that SuperiorGAT consistently achieves lower reconstruction error and improved geometric consistency compared to traditional interpolation methods, PointNet-based models, and deeper GAT baselines. These results indicate that targeted architectural refinement is more critical than increasing graph depth for accurate reconstruction in scenarios characterized by structured vertical sparsity.
Qualitative analyses based on X–Z projection profiles further confirm that SuperiorGAT preserves structural integrity across diverse scenes with minimal vertical distortion. Since the proposed method operates independently of temporal or RGB data and offers a significant reduction in latency compared to multi-layer graph architectures, it provides a practical and scalable framework for autonomous perception. Future work will focus on optimizing the neighbor-search mechanisms to further reduce inference time and extending the framework to dynamic dropout patterns to improve stability in highly transient environments.

### v3

@@ Introduction
Figure 5: Sensitivity analysis of the neighborhood size $k$ showing (a) the trade-off between vertical reconstruction error ($RMSE_{z}$) and inference latency, and (b) the corresponding geometric fidelity measured by Chamfer distance.
As shown in Fig. 5(a), increasing the neighborhood size from $k=5$ to $k=10$ leads to a significant reduction in vertical reconstruction error, achieving the lowest $RMSE_{z}$ of 0.1808 m at $k=10$. Beyond this point, further increasing $k$ results in a gradual increase in $RMSE_{z}$, indicating that excessive neighborhood aggregation introduces over-smoothing effects that degrade local vertical precision.
In contrast, Fig. 5(b) shows that the Chamfer distance continues to decrease as $k$ increases, reflecting improved global geometric alignment when larger spatial contexts are considered. This divergence between vertical accuracy and global shape fidelity highlights an inherent trade-off: smaller neighborhoods preserve fine-grained elevation details, while larger neighborhoods favor smoother but less locally precise reconstructions.
From a computational perspective, inference latency increases approximately linearly with $k$, rising from 7.08 ms at $k=5$ to 12.09 ms at $k=20$. Considering reconstruction accuracy, geometric fidelity, and runtime jointly, $k=10$ represents a clear efficiency–accuracy balance point. Consequently, $k=10$ is selected as the default configuration and is used in all subsequent experiments.
Cross-Environment Evaluation on KITTI
After fixing the neighborhood size to $k=10$, we evaluate the generalization capability of SuperiorGAT across multiple real-world operating conditions using five distinct KITTI sequences: Residential, City, Person, Campus, and Road. These environments exhibit substantial variability in scene structure, object density, and terrain complexity, providing a rigorous test of cross-domain robustness.
Quantitative results are summarized in Table 1. SuperiorGAT consistently achieves strong reconstruction performance across all environments, maintaining an $RMSE_{xyz}$ below 0.11 m in every sequence while simultaneously producing the lowest or near-lowest Chamfer distances among learning-based methods. This indicates that the proposed model effectively preserves both vertical accuracy and global geometric structure across diverse scene types.
In contrast, Nearest Neighbor interpolation occasionally yields competitive $RMSE_{z}$ values in specific environments; however, this apparent advantage does not translate to improved three-dimensional fidelity. The corresponding Chamfer distances are substantially higher, reflecting poor global alignment and the presence of structural artifacts. This behavior highlights the limitation of purely local interpolation schemes, which fail to model spatial context beyond immediate neighbors.
Compared to the standard GAT baseline, SuperiorGAT demonstrates improved reconstruction accuracy while reducing inference latency from approximately 13 ms to 8.6 ms per frame. Importantly, this latency remains stable across all evaluated environments, indicating that the proposed architecture does not rely on environment-specific tuning. Overall, these results confirm that SuperiorGAT generalizes reliably across heterogeneous KITTI scenes while maintaining inference latency compatible with real-time autonomous perception pipelines. This efficiency gain is achieved by avoiding stacked attention layers and instead employing a single beam-aware graph attention layer augmented with gated residual fusion and a lightweight feed-forward refinement block, which reduces computational overhead without sacrificing reconstruction fidelity.
@@
Dataset Method $RMSE_{z}$ (m) $RMSE_{xyz}$ (m) Chamfer Distance (m) Time (ms)
@@
Robustness to Structured Beam Dropout
To assess robustness under realistic sensor degradation, we evaluate SuperiorGAT under structured beam dropout on both the KITTI and nuScenes datasets. Unlike random point removal, structured beam dropout eliminates contiguous vertical scan lines, emulating reduced-resolution LiDAR sensors, partial sensor failure, or adverse environmental conditions. Dropout rates of 12.5%, 25%, and 37.5% are considered to progressively increase reconstruction difficulty.
Figs. 6a and 6b illustrate the effect of increasing beam dropout on vertical reconstruction accuracy for KITTI and nuScenes, respectively. Across both datasets, SuperiorGAT consistently achieves the lowest $RMSE_{z}$ at all dropout levels and exhibits a significantly slower degradation rate compared to baseline methods. This behavior indicates that the proposed attention-based architecture effectively leverages remaining spatial context to recover missing vertical structure even under severe information loss.
(a) KITTI dataset
(b) nuScenes dataset
Figure 6: Vertical reconstruction error ($RMSE_{z}$) as a function of structured beam dropout rate for (a) KITTI and (b) nuScenes. SuperiorGAT exhibits consistently lower error and more graceful degradation compared to baseline methods across increasing dropout levels.
Quantitative results are further detailed in Tables 2 and 3. On KITTI, SuperiorGAT maintains stable performance as dropout increases, with only moderate growth in $RMSE_{z}$ and Chamfer distance, while Linear Interpolation shows large and largely invariant errors regardless of dropout severity. The GAT baseline degrades more rapidly, highlighting its limited ability to adapt to structured beam removal.
A similar trend is observed on nuScenes, which operates with a lower vertical resolution 32-beam LiDAR sensor. Despite the inherently more challenging reconstruction setting, SuperiorGAT preserves a clear performance margin over both interpolation-based and graph-based baselines across all dropout rates. Notably, inference time remains nearly constant for all dropout configurations on both datasets, confirming that robustness is achieved through learned spatial reasoning rather than increased computational cost.
Overall, these results demonstrate that SuperiorGAT degrades gracefully under structured beam dropout and generalizes robustly across datasets with different LiDAR configurations, reinforcing its suitability for deployment in real-world autonomous perception scenarios.
Table 2: Robustness under structured beam dropout on KITTI (mean over frames, averaged over three runs).
Dropout Method RMSEZ (m) $\downarrow$ Chamfer (m) $\downarrow$ Normal Consistency $\uparrow$ Inference Time (s) $\downarrow$
12.5% Linear Interp 0.50 0.20 0.81 0.03
GAT Baseline 0.21 0.17 0.89 0.04
SuperiorGAT (Ours) 0.12 0.15 0.88 0.02
25% Linear Interp 0.47 0.20 0.80 0.03
GAT Baseline 0.31 0.24 0.85 0.04
SuperiorGAT (Ours) 0.20 0.15 0.85 0.02
37.5% Linear Interp 0.53 0.21 0.78 0.03
GAT Baseline 0.26 0.26 0.83 0.04
SuperiorGAT (Ours) 0.16 0.16 0.84 0.02
Table 3: Robustness under structured beam dropout on nuScenes (mean over frames, averaged over three runs).
Dropout Method RMSEZ (m) $\downarrow$ Chamfer (m) $\downarrow$ Normal Consistency $\uparrow$ Inference Time (s) $\downarrow$
12.5% Linear Interp 16.83 9.15 0.49 0.03
GAT Baseline 5.27 2.63 0.68 0.04
SuperiorGAT (Ours) 3.32 1.37 0.75 0.02
25% Linear Interp 16.84 9.15 0.49 0.03
GAT Baseline 5.30 2.65 0.68 0.04
SuperiorGAT (Ours) 3.35 1.40 0.74 0.02
37.5% Linear Interp 17.25 9.32 0.61 0.03
GAT Baseline 5.60 2.90 0.64 0.04
SuperiorGAT (Ours) 3.75 1.55 0.72 0.02
Qualitative Reconstruction Analysis
Figure 7 presents a qualitative comparison between the original ground-truth LiDAR data and the reconstructed vertical profiles generated by SuperiorGAT across the Person, Road, Campus, and City datasets. For each scene, a camera-view reference is shown on the left to provide operational context, while the corresponding 2D X–Z projection is displayed on the right, illustrating the reconstructed vertical structure as a function of horizontal distance.
For visual clarity, the reconstructed point clouds are displayed using a sampling stride of 15. Despite this downsampling, SuperiorGAT preserves high geometric fidelity, accurately recovering vertical structural height ($Z$) while maintaining smooth and coherent profiles along the horizontal axis ($X$). Across all evaluated environments, the reconstructed profiles closely follow the ground-truth geometry, capturing elevation changes, slopes, and terrain undulations without introducing spurious discontinuities.
In contrast to interpolation-based methods, which typically produce staircase artifacts and fragmented vertical profiles, SuperiorGAT maintains consistent vertical alignment and continuity. These visual results further confirm the model’s ability to integrate local and global spatial context, complementing the quantitative improvements reported in Tables 1–3 and the robustness trends observed in Figs. 6a and 6b.
Overall, the qualitative comparisons reinforce the numerical findings and demonstrate that SuperiorGAT achieves robust and geometrically consistent vertical reconstruction across diverse environments and structured beam dropout conditions.
@@ Conclusion
This paper presented SuperiorGAT, a graph attention–based framework for reconstructing missing elevation information in sparse LiDAR point clouds under structured beam dropout. By representing LiDAR scans as beam-aware graphs and augmenting standard graph attention networks with gated residual fusion and a lightweight feed-forward refinement module, the proposed approach effectively recovers lost vertical geometry while maintaining computational efficiency without increasing network depth.
Extensive experiments across multiple KITTI environments, including Person, Road, Campus, and City, demonstrate that SuperiorGAT consistently achieves lower reconstruction error and improved geometric consistency compared to traditional interpolation techniques, PointNet-based models, and deeper GAT baselines. Robustness evaluations under increasing structured beam dropout further show that SuperiorGAT degrades gracefully and preserves a clear performance margin across both KITTI and nuScenes datasets, highlighting its ability to generalize across different LiDAR configurations and sensing resolutions.
Qualitative analyses based on X–Z projection profiles corroborate the quantitative findings, confirming that SuperiorGAT preserves structural continuity and vertical alignment across diverse scenes with minimal distortion. As the proposed framework operates solely on geometric information and does not rely on temporal aggregation or RGB data, it offers a practical and scalable solution for LiDAR reconstruction in autonomous perception pipelines. Future work will explore extensions to dynamic dropout patterns and temporal integration to further enhance reconstruction stability in highly dynamic environments. Future work will investigate component-level ablation and cross-dataset generalization to further characterize the model’s robustness under diverse sensor configurations and dynamic dropout patterns.